Solve for the unknown number of years in each of the following (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.: Interest Rate Future Value 8% $ 1,496 2.496 392,101 403,794 Present Value Years $ 630 880 19,100 22,200 12 18 1A
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

a.

1496=630*(1.08)^n

(1496/630)=(1.08)^n

Taking log on both sides;

log (1496/630)=n*log (1.08)

n=log(1496/630)/log 1.08

=11.24 years(Approx).

b.

2496=880*(1.12)^n

(2496/880)=(1.12)^n

Taking log on both sides;

log (2496/880)=n*log (1.12)

n=log(2496/880)/log 1.12

=9.20 years(Approx).

c.

392101=19100*(1.18)^n

(392101/19100)=(1.18)^n

Taking log on both sides;

log (392101/19100)=n*log (1.18)

n=log(392101/19100)/log 1.18

=18.26 years(Approx).

d.

403794=22200*(1.14)^n

(403794/22200)=(1.14)^n

Taking log on both sides;

log (403794/22200)=n*log (1.14)

n=log(403794/22200)/log 1.14

=22.14 years(Approx).